OpenStudy (anonymous):

Which is the prime factorization of 2,240? A. 2^3 • 8 • 5 • 7 B. 2^6 • 15 • 7 C. 2^6 • 5 • 7 D. 2^5 • 5 • 7

OpenStudy (anonymous):

I think A would be the correct answer.. but can anyone explain how to find it? the prime factorization?

OpenStudy (goldphenoix):

No. A is not correct.

OpenStudy (anonymous):

correct answer is C

OpenStudy (goldphenoix):

8 is not a prime number. :|

OpenStudy (anonymous):

ok then can you explain...how to find it

OpenStudy (goldphenoix):

Also for b, 15 is not a prime number. So it's either C or D.

OpenStudy (anonymous):

I think it is C!

OpenStudy (anonymous):

C

OpenStudy (skullpatrol):

The prime factorization of an integer is found by using the primes in order as divisors, dividing by each as many times as possible.

OpenStudy (anonymous):

you can dismiss the first two answers because they have 8 and 15 (not a prime numbers) now just do the math: 2^6 * 5 * 7 = 2240 2^5 * 5 * 7 = 1120 to find a prime factorization just divide the number with prime numbers, starting from the smallest: 2240/2=1140 1140/2=560 560/2=280 280/2=140 140/2=70 70/2=35 you can't divide with 2 anymore, take the next prime number, which is 3 but, unfortunately, you can't divide 35 with 3. proceed with the next prime number, which is 5. 35/5=7 can't divide with 5 anymore, proceed to the next prime number, 7 7/7 = 1 and now just product all the numbers you used to divide the starting number: 2*2*2*2*2*2*5*7 that is prime factorization

OpenStudy (goldphenoix):



OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

euclidian algorithm is always fun...

OpenStudy (goldphenoix):

You divide the number 2240 by the lowest prime number, 2. Keep dividing until the number cannot be divisible by 2, such as 35. Then you divide it by the second lowest, 3. But you cannot divide 35 into 3. So you go the third lowest, 5. 5 is divisible by 35. 35 divided by 5 is 7. 7 is a prime number, so you count that to. You then count all of the number that you divided with. You would end up with: 2^6 • 5 • 7

OpenStudy (skullpatrol):

Try the primes in order 2, 3, 5,... as divisors. Divide by each prime as many times as possible before going on to the next prime.

Directrix (directrix):

@RH If you are allowed to use technology in your work, look at the online prime factors calculator at the following site: http://www.calculatorsoup.com/calculators/math/prime-factors.php

OpenStudy (anonymous):

Hi, You can use fraction calculator to calculate the factorization of 2,240. http://www.acalculator.com/fraction-operations-mixed-numbers-calculator.html

OpenStudy (skullpatrol):

Welcome back @Directrix :-)